# XSNAP: An X-ray Supernova Analysis Pipeline with Application to the Type II Supernova 2024ggi

### v1

@@ Abstract
X-ray observations of Type II supernovae (SNe II) probe the physics of supernova (SN) shocks and the mass-loss histories of their progenitor stars. We present multi-epoch, X-ray observations of SN II 2024ggi ($D\approx 7.2$ Mpc) from Swift-XRT, Chandra X-ray Observatory and XMM-Newton, which cover $\sim 1-344$ days since first light. We analyze these observations using a new open-source Python package called XSNAP, which standardizes a unified command-line interface for instrument-specific reduction and spectral extraction. XSNAP introduces application programming interfaces for per-epoch spectral modeling through PyXspec and emcee Markov chain Monte Carlo fitting. We employ XSNAP to model the multi-epoch X-ray spectra of SN 2024ggi with an absorbed thermal bremsstrahlung model and calculate a steady progenitor mass-loss rate of $(6.2\pm 0.2)\times 10^{-5}\,M_{\odot}\,\mathrm{yr^{-1}}$ $(v_{\rm wind}=20\ \rm km\ s^{-1})$, for which the detected X-ray emission traces the final 117 years before explosion. The software is publicly available on GitHub, with a released package on the Python Package Index (PyPI).
@@ 1 Introduction
An example of a well-studied Type IIP SN in the X-rays is SN 2023ixf, discovered by K. Itagaki on 2023 May 19 in the M101 galaxy (K. Itagaki, 2023). It was first observed in at X-ray wavelengths with the Swift X-ray Telescope (XRT) at $\sim 1$ day after first light, which was then complimented with follow-up observations from the Chandra X-ray Observatory (CXO), XMM-Newton, and Nuclear Spectroscopic Telescope Array (NuSTAR) (W. Jacobson-Galán, 2025). Multiple analyses have been done with the absorbed plasma model (P. Chandra et al., 2024a) and the absorbed Bremsstrahlung model (B. W. Grefenstette et al., 2023; E. A. Zimmerman et al., 2024; S. Panjkov et al., 2024; A. J. Nayana et al., 2025; W. V. Jacobson-Galán et al., 2025a) across the soft and hard X-rays, with the X-ray luminosity of $\sim 10^{39}-10^{40}\ \rm erg\ s^{-1}$ within the first $\sim 10^{2}\ \rm days$. The findings of these studies are consistent and imply that SN 2023ixf interacted with dense, confined CSM and the red supergiant (RSG) progenitor mass-loss rate was around $10^{-4}-10^{-3}\ \rm M_{\odot}\ yr^{-1}$, one to two order magnitude higher than the upper-limit of SNe IIP detected in X-rays (V. V. Dwarkadas, 2025).
@@
SN 2024ggi, another nearby Type II SN (W. Hoogendam et al., 2024; Q. Zhai et al., 2024), is a well-monitored SN across multiple wavelengths: ($\gamma$-rays (G. Marti-Devesa & Fermi-LAT Collaboration, 2024), X-rays (J. Zhang et al., 2024a; R. Margutti & B. Grefenstette, 2024), optical bands (T. W. Chen et al., 2024; T. Killestein et al., 2024; B. Kumar et al., 2024; F. D. Romanov, 2024; L. Wyrzykowski et al., 2025), radio frequencies (S. Ryder et al., 2024), and centimeter wavelengths (P. Chandra et al., 2024b)). The SN was first discovered by the Asteroid Terrestrial-impact Last Alert System (ATLAS) on 2024-04-11 (MJD 60411.14) (J. Tonry et al., 2024; S. Srivastav et al., 2024; T. W. Chen et al., 2025) and was officially classified as a Type IIP (J. Zhang et al., 2024b; T. W. Chen et al., 2025). Previous studies have analyzed UV/optical/IR observations and inferred that the RSG progenitor mass-loss rate was $10^{-3}-10^{-2}\ \rm M_{\odot}\ yr^{-1}$ (M. Shrestha et al., 2024; J. Zhang et al., 2024b; W. V. Jacobson-Galán et al., 2024a; A. Aryan et al., 2025; T. W. Chen et al., 2025; K. Ertini et al., 2025). Additionally, D. Xiang et al. (2024) estimated the progenitor mass-loss rate is $3\times 10^{-6}\ \rm M_{\odot}\ yr^{-1}$ from pre-explosion imaging and suggested that SN 2024ggi had an enhancement of mass-loss rate in the century leading up to the explosion. Use of nebular spectroscopy has indicated a progenitor mass in the range of $11-15.2\ \rm M_{\odot}$ (D. Xiang et al., 2024; A. Aryan et al., 2025; L. Dessart et al., 2025; K. Ertini et al., 2025; L. Ferrari et al., 2025; E. Hueichapán et al., 2025). In this work, we present and analyze the X-ray observations of SN 2024ggi from Chandra X-ray Observatory, Swift X-ray Telescope, and XMM-Newton. SN 2024ggi is located at $(\alpha=\rm 11^{h}18^{m}22.09^{s},\delta=-32^{\circ}50^{\prime}15.26^{\prime\prime})$ in the host galaxy NGC 3621. We adopt a time of first light corresponding to MJD $60410.80\pm 0.34$ days, redshift of $z=0.002215$ (W. V. Jacobson-Galán et al., 2024a), and a distance of $7.2\pm 0.2$ Mpc (A. Saha et al., 2006).
@@
We present multi-epoch X-ray observations of SN 2024ggi in Section 2. Our analysis of SN 2024ggi and its surrounding CSM are described in Section 3. In Section 4, we introduce the pipeline we have developed, XSNAP, and its design and availability. We discuss our findings in Section 5 and summarize them in Section 6.
2 Observations
@@
(days) $\left(10^{22}\ \rm cm^{-2}\right)$ (keV) (Absorbed) (Unabsorbed) ($10^{-4}$)
@@
Figure 2: Left: Plot of fitted unabsorbed $0.3-10\ \rm keV$ X-ray luminosity light curve from XMM-Newton (red circles), CXO (cyan squares), and Swift-XRT (orange triangles) observations. The red dashed line is the fitted line with $L\propto t^{-0.99}$. Right: Plot of unabsorbed $0.3-10\ \rm keV$ X-ray luminosity light curve of SN 2024ggi compared to a sample of Type IIP SNe. References: E. M. Schlegel (1999, 2001), D. Pooley et al. (2002), D. Pooley & W. H. G. Lewin (2002, 2004), K. Misra et al. (2007), S. Immler et al. (2007), S. Immler & P. J. Brown (2012), S. Chakraborti et al. (2013, 2016), T. Szalai et al. (2019), I. Irani et al. (2024), A. J. Nayana et al. (2025)
@@ 2.1 Chandra X-ray Observatory
The Chandra X-ray Observatory (CXO) ACIS-S observed SN 2024ggi in two epochs: one on 2024 April 21 (ObsID: 29383; $\delta t=10.90\ \rm days$; PI: E. Zimmerman) with an exposure time of 14.88 ks and the other on 2024 April 26 (ObsID: 29384; $\delta t=16.20\ \rm days$; PI: E. Zimmerman) with an exposure time of 14.58 ks. Details of these two observations are reported in Table 3.
@@
We reduced both observations using CIAO v4.17 (A. Fruscione et al., 2006) and their corresponding calibration files. Source detection and estimating source-free background regions were utilized through ximage v4.5.1 from HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014)Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), Heasarc) with signal-to-noise ratio threshold of three $(\rm SNR_{thresh}=3)$ and human vetting with SAOImage DS9 (W. A. Joye & E. Mandel, 2003). For each epoch, we extracted the spectrum via specextract and used a circular source region with a radius of $2^{\prime\prime}$ and circular background region with a radius of $45^{\prime\prime}$.
2.2 XMM-Newton
@@
Similar to CXO, XMM-Newton (XMM) observed SN 2024ggi in two epochs: one on 2024 June 04 (ObsID: 0882480901; $\delta t=55.03\ \rm days$; PI: W. V. Jacobson-Gálan) and the other on 2024 July 5 (ObsID: 0882481001; $\delta t=85.40\ \rm days$; PI: W. V. Jacobson-Gálan). Details of these two observations are reported in Table 3.
@@
We reduced both observations from the three instruments of XMM, i.e. European Photon Imaging Cameras (EPIC)-pn, MOS1, and MOS2, using the Scientific Analysis System (SAS) v22.1.0 (SAS development team, 2014) and corresponding calibration files. Source detection and estimating source-free background regions were utilized through ximage v4.5.1 from HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014)Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), Heasarc) with signal-to-noise ratio threshold of three $(\rm SNR_{thresh}=3)$ and human vetting with SAOImage DS9 (W. A. Joye & E. Mandel, 2003). For each epoch, we extracted the spectrum via evselect and used a circular source region with a radius of $25^{\prime\prime}$ and circular background region with a radius of $125^{\prime\prime}$.
2.3 Swift X-ray Telescope
@@
The X-Ray Telescope (XRT; D. N. Burrows et al. 2005) onboard the Neil Gehrels Swift Observatory (N. Gehrels et al., 2004) started observing SN 2024ggi on 2024 April 11 UT 14:04:10 $(\delta t=0.79\ \rm days)$. In this work, we analyzed XRT datasets collected until 2025 March 20. As each observation has relatively small exposure time $(\sim 1.5\ \rm ks)$, we binned a few epochs together, ranging from 10-day to 30-day intervals until the source was detected. However, SN 2024ggi was not detected in XRT images after 2024 June 02 $(\delta t=52.20\ \rm days)$ and these observations were analyzed as upper-limits (see Section 3). Details of each stacked observation is reported in Table 3.
@@
We reduced each stacked observation using HEASoft v6.35 (Nasa High Energy Astrophysics Science Archive Research Center (2014)Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), Heasarc), the Swift XRT Data Analysis Software (SWXRTDAS; version 3.7.0), and the Swift Calibration Database (CALDB) version 20250609. Source detection and estimating source-free background regions were utilized through ximage v4.5.1 from HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014)Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), Heasarc) with signal-to-noise ratio threshold of three $(\rm SNR_{thresh}=3)$ and human vetting with SAOImage DS9 (W. A. Joye & E. Mandel, 2003). For each epoch, we extracted the spectrum via xrtpipeline, followed the standard practice from P. A. Evans et al. (2009); R. Margutti et al. (2012), and used a circular source region with a radius of $25^{\prime\prime}$ and circular background region with a radius of $125^{\prime\prime}$.
3 Analysis
3.1 X-ray Spectral Modeling
@@
We modeled and fit each observation in Table 3 with PyXspec (C. Gordon & K. Arnaud, 2021), the Python interface to XSPEC (K. A. Arnaud, 1996). We fit them separately except for XMM-Newton data where we did a joint fit for all three instruments, EPIC-pn, MOS1, and MOS2. In this work, we adopt the solar abundances from M. Asplund et al. (2009), i.e., abund aspl within PyXspec. We set the Galactic neutral hydrogen (NH) column density in the direction of SN2024ggi of $\mathrm{NH}_{\mathrm{MW}}=6.59\cdot 10^{20}\ \mathrm{cm}^{-2}$ (HI4PI Collaboration et al., 2016). For the spectral fits, we restricted the Chandra/ACIS data to the 0.58keV0.5-8\ $\mathrm${keV} energy range, while the Swift-XRT and XMM-Newton spectra were fit in the 0.310keV0.3-10\ $\mathrm${keV} band.
@@
We then fit these data using an absorbed thermal Bremsstrahlung model (tbabs*ztbabs*bremss), but were unable to constrain the model temperature. Therefore, we estimated and fixed the temperatures of each observation using the self-similar solution $T=34\ \rm keV\ (t/13d)^{-0.25}$ (P. Chandra et al., 2024a). This temperature evolution follows the self-similar solutions by R. A. Chevalier (1982b); R. A. Chevalier & C. Fransson (2017) for a standard wind with a constant mass-loss rate, $\rho\propto r^{-2}$. For such a case, the post-shock temperature evolves as $T\propto t^{-2/(n-2)}$, which becomes $T\propto t^{-0.25}$ after assuming $n=10$. We adopted a normalization of $T=34~\rm keV$ at $t=13~\rm d$, as utilized for SN 2023ixf by P. Chandra et al. (2024a). Finally, we successfully modeled the intrinsic NH and Bremsstrahlung normalization and estimated the absorbed and unabsorbed fluxes in $0.3-10\ \rm keV$. We also repeat this analysis while adopting the temperature evolution of SN 2023ixf, $T\propto t^{-0.5}$ (A. J. Nayana et al., 2025), which yields consistent results to those found with a self-similar solution temperature decline rate. Based on the resulting best‐fit parameters, the X-ray spectra of SN 2024ggi reveal SN ejecta interaction with dense CSM: (i) the emission was highly absorbed at early times, (ii) the absorption decreased with time, and (iii) of thermal nature, consistent with previous studies (R. Margutti et al., 2017; B. P. Thomas et al., 2022). The best-fit parameters are reported in Table 1 and contour plots of the first four epochs data from Swift-XRT and CXO are shown in 4. In addition, the fitted spectrum for CXO and XMM-Newton observations are shown in Figure 1.
@@
For observations where SN 2024ggi was not detected, we simulated the spectrum using the FakeIt command from PyXspec and used the faked spectrum flux and count rate ratio to estimate the flux upper-limit in $0.3-10\ \rm keV$. We set the intrinsic NH for non-detections to be $\rm NH_{int}=0$ and did not include their best-fit parameters for further analysis. We converted all the unabsorbed fluxes to luminosities and fit a power-law with emcee Markov chain Monte Carlo (MCMC). We found that the intrinsic $0.3-10\ \rm keV$ luminosity has a decay of $L_{X}\propto t^{-0.99}$ at $\delta t\gtrsim 8.36$ days. Additionally, in Figure 2, we compare the X-ray light curve to a sample of SNe II compiled by I. Irani et al. (2024); A. J. Nayana et al. (2025).
3.2 CSM Density Profile Analysis

### v2

@@ Abstract
X-ray observations of Type II supernovae (SNe II) probe the physics of supernova (SN) shocks and the mass-loss histories of their progenitor stars. We present multi-epoch, X-ray observations of SN II 2024ggi ($D\approx 7.2$ Mpc) from Swift-XRT, Chandra X-ray Observatory and XMM-Newton, which cover $\sim 1-344$ days since first light. We analyze these observations using a new open-source Python package called XSNAP, which standardizes a unified command-line interface for instrument-specific reduction and spectral extraction. XSNAP introduces application programming interfaces for per-epoch spectral modeling through PyXspec and emcee Markov chain Monte Carlo fitting. We employ XSNAP to model the multi-epoch X-ray spectra of SN 2024ggi with an absorbed thermal bremsstrahlung model and calculate a steady progenitor mass-loss rate of $(6.2\pm 0.2)\times 10^{-5}\,M_{\odot}\,\mathrm{yr^{-1}}$ $(v_{\rm wind}=20\ \rm km\ s^{-1})$, for which the detected X-ray emission traces the final $\sim 117$ years before explosion. The software is publicly available on GitHub, with a released package on the Python Package Index (PyPI).
@@ I Introduction
An example of a well-studied Type IIP SN in the X-rays is SN 2023ixf, discovered by K. Itagaki on 2023 May 19 in the M101 galaxy (K. Itagaki, 2023). It was first observed in at X-ray wavelengths with the Neil Gehrels Swift Observatory X-ray Telescope (Swift-XRT) at $\sim 1$ day after first light, which was then complimented with follow-up observations from the Chandra X-ray Observatory (CXO), XMM-Newton (XMM), and Nuclear Spectroscopic Telescope Array (NuSTAR) (W. Jacobson-Galán, 2025). Multiple analyses have been done with the absorbed plasma model (P. Chandra et al., 2024a) and the absorbed Bremsstrahlung model (B. W. Grefenstette et al., 2023; E. A. Zimmerman et al., 2024; S. Panjkov et al., 2024; A. J. Nayana et al., 2025; W. V. Jacobson-Galán et al., 2025a) across the soft and hard X-rays, with the X-ray luminosity of $\sim 10^{39}-10^{40}\ \rm erg\ s^{-1}$ within the first $\sim 10^{2}\ \rm days$. The findings of these studies are consistent and imply that SN 2023ixf interacted with dense, confined CSM and the red supergiant (RSG) progenitor mass-loss rate was around $10^{-4}-10^{-3}\ \rm M_{\odot}\ yr^{-1}$, one to two order magnitude higher than the upper-limit of SNe IIP detected in X-rays (V. V. Dwarkadas, 2025).
@@
SN 2024ggi, another nearby Type II SN (W. Hoogendam et al., 2024; Q. Zhai et al., 2024), is a well-monitored SN across multiple wavelengths: ($\gamma$-rays (G. Marti-Devesa & Fermi-LAT Collaboration, 2024), X-rays (J. Zhang et al., 2024a; R. Margutti & B. Grefenstette, 2024), optical bands (T. W. Chen et al., 2024; T. Killestein et al., 2024; B. Kumar et al., 2024; F. D. Romanov, 2024; L. Wyrzykowski et al., 2025), radio frequencies (S. Ryder et al., 2024), and centimeter wavelengths (P. Chandra et al., 2024b)). The SN was first discovered by the Asteroid Terrestrial-impact Last Alert System (ATLAS) on 2024-04-11 (MJD 60411.14) (J. Tonry et al., 2024; S. Srivastav et al., 2024; T. W. Chen et al., 2025) and was officially classified as a Type IIP (J. Zhang et al., 2024b; T. W. Chen et al., 2025). Previous studies have analyzed UV/optical/IR observations and inferred that the RSG progenitor mass-loss rate was $10^{-3}-10^{-2}\ \rm M_{\odot}\ yr^{-1}$ (M. Shrestha et al., 2024; J. Zhang et al., 2024b; W. V. Jacobson-Galán et al., 2024a; A. Aryan et al., 2025; T. W. Chen et al., 2025; K. Ertini et al., 2025). Additionally, D. Xiang et al. (2024) estimated the progenitor mass-loss rate is $3\times 10^{-6}\ \rm M_{\odot}\ yr^{-1}$ from pre-explosion imaging and suggested that SN 2024ggi had an enhancement of mass-loss rate in the century leading up to the explosion. Use of nebular spectroscopy has indicated a progenitor mass in the range of $11-15.2\ \rm M_{\odot}$ (D. Xiang et al., 2024; A. Aryan et al., 2025; L. Dessart et al., 2025; K. Ertini et al., 2025; L. Ferrari et al., 2025; E. Hueichapán et al., 2025). In this work, we present and analyze the X-ray observations of SN 2024ggi from CXO, Swift-XRT, and XMM. SN 2024ggi is located at $(\alpha=\rm 11^{h}18^{m}22.09^{s},\delta=-32^{\circ}50^{\prime}15.26^{\prime\prime})$ in the host galaxy NGC 3621. We adopt a time of first light corresponding to MJD $60410.80\pm 0.34$ days, redshift of $z=0.002215$ (W. V. Jacobson-Galán et al., 2024a), and a distance of $7.2\pm 0.2$ Mpc (A. Saha et al., 2006).
@@
In Section II, we describe the design, availability and applications of the “X-ray Supernova Analysis Pipeline (XSNAP).” In Section III, we present multi-epoch X-ray observations of SN 2024ggi, and our analysis of SN 2024ggi and its surrounding CSM are described in Section IV. We discuss our findings in Section V and summarize them in Section VI.
II The Pipeline
Each X-ray observatory, such as CXO, XMM, and Swift-XRT, and NuSTAR, employs mission-specific data formats, calibration pipelines, and analysis software. Consequently, this hinders reproducibility in analyzing X-ray observations as astronomers need to work out distinct guidelines for each mission. While there have been past efforts to streamline the process, such as YAXX [^1] https://cxc.harvard.edu/contrib/yaxx (Yet Another X-ray Xtractor; T. Aldcroft 2006), which supported both CXO and XMM-Newton data, these tools are no longer actively maintained. More recent softwares, like XGA [^2] https://xga.readthedocs.io/en/latest/ (X-ray: Generate and Analyze; D. J. Turner et al. 2022) or BXA [^3] https://johannesbuchner.github.io/BXA/ (Bayesian X-ray Analysis; J. Buchner et al. 2014), offer modern functionality but are limited to a single instrument (e.g., XMM), not designed as end-to-end pipelines from spectral extraction, or not specifically developed for SNe, even though some may still be adaptable for such studies. At present, there is no open-source, Python-based pipeline that unifies spectral extraction and analysis of SN X-ray spectra from CXO, XMM, Swift-XRT, and NuSTAR. Consequently, we developed XSNAP (X-ray Supernova Analysis Pipeline), an open-source Python package intended to unify end-to-end X-ray data processes from spectral extraction to CSM analysis.
XSNAP is organized into two stages: spectrum extraction and spectral modeling and is designed to streamline all the processes described in Sections III and IV, including source detection, deriving count-rates, spectral fitting and deriving unshocked CSM densities. A typical XSNAP workflow consits of (i) generating source and background regions, (ii) extracting and reducing raw spectrum through the Python Command-line Interface (CLI) scripts, and (iii) analyzing spectrum from the Python Application Programming Interface (API) modules. This unified workflow is applied to the analysis of SN 2024ggi presented in Sections III and IV. The spectrum extraction component standardizes the procedures for observations from CXO, XMM, and Swift-XRT (Section III), implementing them as dedicated Python CLI scripts. In addition, a script is made seperately for observations from NuSTAR; we follow the standard guide from NuSTAR Data Analysis Software (NuSTARDAS [^4] https://heasarc.gsfc.nasa.gov/docs/nustar/analysis/nustar_quickstart_guide.pdf) (F. A. Harrison et al., 2013) inside HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014)Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), Heasarc). XSNAP also provides a script for generating source and background regions, making use of ximage and (optionally) SAOImage DS9 (W. A. Joye & E. Mandel, 2003). Similarly, we developed the spectrum and CSM analysis sector by wrapping all the processes from Section IV to each separate Python API module. We then packaged the pipeline in a Python Package Index (PyPI) environment and published it to Github [^5] https://github.com/fercananything/XSNAP/, PyPI [^6] https://pypi.org/p/xsnap/, and its own documentation website [^7] https://xsnap.org/.
III Observations
@@
(days) $\left(10^{22}\ \rm cm^{-2}\right)$ (keV) (Absorbed) (Unabsorbed) ($10^{-4}\ \rm\mathbf{cm^{-5}}$)
@@
Figure 2: Left: Plot of fitted unabsorbed $0.3-10\ \rm keV$ X-ray luminosity light curve from XMM (red circles), CXO (cyan squares), and Swift-XRT (orange triangles) observations. The red dashed line is the fitted line with $L\propto t^{-0.99}$. Right: Plot of unabsorbed $0.3-10\ \rm keV$ X-ray luminosity light curve of SN 2024ggi compared to a sample of Type IIP SNe. References: E. M. Schlegel (1999, 2001), D. Pooley et al. (2002), D. Pooley & W. H. G. Lewin (2002, 2004), K. Misra et al. (2007), S. Immler et al. (2007), S. Immler & P. J. Brown (2012), S. Chakraborti et al. (2013, 2016), T. Szalai et al. (2019), I. Irani et al. (2024), A. J. Nayana et al. (2025)
@@ III.1 Chandra X-ray Observatory
The CXO ACIS-S observed SN 2024ggi in two epochs: one on 2024 April 21 (ObsID: 29383; $\delta t=10.90\ \rm days$; PI: E. Zimmerman) with an exposure time of 14.88 ks and the other on 2024 April 26 (ObsID: 29384; $\delta t=16.20\ \rm days$; PI: E. Zimmerman) with an exposure time of 14.58 ks. Details of these two observations are reported in Table 2.
@@
We reduced both observations using CIAO v4.17 (A. Fruscione et al., 2006) and their corresponding calibration files. Source detection and estimating source-free background regions were utilized through ximage v4.5.1 from HEASoft with signal-to-noise ratio threshold of three $(\rm SNR_{thresh}=3)$ and human vetting with SAOImage DS9. For each epoch, we extracted the spectrum via specextract and used a circular source region with a radius of $2^{\prime\prime}$ and circular background region with a radius of $45^{\prime\prime}$.
III.2 XMM-Newton
@@
Similar to CXO, XMM observed SN 2024ggi in two epochs: one on 2024 June 04 (ObsID: 0882480901; $\delta t=55.03\ \rm days$; PI: W. V. Jacobson-Gálan) and the other on 2024 July 05 (ObsID: 0882481001; $\delta t=85.40\ \rm days$; PI: W. V. Jacobson-Gálan). Details of these two observations are reported in Table 2.
@@
We reduced both observations from the three instruments of XMM, i.e. European Photon Imaging Cameras (EPIC)-pn, MOS1, and MOS2, using the Scientific Analysis System (SAS) v22.1.0 (SAS development team, 2014) and corresponding calibration files. Source detection and estimating source-free background regions were utilized through ximage v4.5.1 from HEASoft with signal-to-noise ratio threshold of three $(\rm SNR_{thresh}=3)$ and human vetting with SAOImage DS9. For each epoch, we extracted the spectrum via evselect and used a circular source region with a radius of $25^{\prime\prime}$ and circular background region with a radius of $125^{\prime\prime}$.
III.3 Swift X-ray Telescope
@@
The Swift-XRT (D. N. Burrows et al., 2005) onboard the Neil Gehrels Swift Observatory (N. Gehrels et al., 2004) started observing SN 2024ggi on 2024 April 11 UT 14:04:10 $(\delta t=0.79\ \rm days)$. In this work, we analyzed Swift-XRT datasets collected until 2025 March 20. As each observation has relatively small exposure time $(\sim 1.5\ \rm ks)$, we binned a few epochs together, ranging from 10-day to 30-day intervals until the source was detected. However, SN 2024ggi was not detected in Swift-XRT images after 2024 June 02 $(\delta t=52.20\ \rm days)$ and these observations were analyzed as upper-limits (see Section IV). Details of each stacked observation is reported in Table 2.
@@
We reduced each stacked observation using HEASoft v6.35, the Swift XRT Data Analysis Software (SWXRTDAS; version 3.7.0), and the Swift Calibration Database (CALDB) version 20250609. Source detection and estimating source-free background regions were utilized through ximage v4.5.1 from HEASoft with signal-to-noise ratio threshold of three $(\rm SNR_{thresh}=3)$ and human vetting with SAOImage DS9. For each epoch, we extracted the spectrum via xrtpipeline, followed the standard practice from P. A. Evans et al. (2009); R. Margutti et al. (2012), and used a circular source region with a radius of $25^{\prime\prime}$ and circular background region with a radius of $125^{\prime\prime}$.
IV Analysis
IV.1 X-ray Spectral Modeling
@@
We modeled and fit each observation in Table 2 with PyXspec (C. Gordon & K. Arnaud, 2021), the Python interface to XSPEC (K. A. Arnaud, 1996). We fit them separately except for XMM data where we did a joint fit for all three instruments, EPIC-pn, MOS1, and MOS2. In this work, we adopted the solar abundances from M. Asplund et al. (2009), i.e., abund aspl within PyXspec. We set the Galactic neutral hydrogen ($\rm N_{H}$) column density in the direction of SN2024ggi of $\mathrm{N}_{\mathrm{H,\ MW}}=6.59\cdot 10^{20}\ \mathrm{cm}^{-2}$ (HI4PI Collaboration et al., 2016). For the spectral fits, we restricted the Chandra/ACIS data to the 18keV1-8\ $\mathrm${keV} energy range, while the Swift-XRT and XMM spectra were fit in the 0.310keV0.3-10\ $\mathrm${keV} band.
@@
We then fit these data using an absorbed thermal Bremsstrahlung model (tbabs*ztbabs*bremss), but were unable to constrain the model temperature. Therefore, we estimated and fixed the temperatures of each observation using the self-similar solution $T=34\ \rm keV\ (t/13d)^{-0.25}$ (P. Chandra et al., 2024a). This temperature evolution follows the self-similar solutions by R. A. Chevalier (1982b); R. A. Chevalier & C. Fransson (2017) for a standard wind with a constant mass-loss rate, $\rho\propto r^{-2}$. For such a case, the post-shock temperature evolves as $T\propto t^{-2/(n-2)}$, which becomes $T\propto t^{-0.25}$ after assuming $n=10$. We adopted a normalization of $T=34~\rm keV$ at $t=13~\rm d$, as utilized for SN 2023ixf by P. Chandra et al. (2024a). Finally, we modeled the intrinsic $\rm N_{H}$ and Bremsstrahlung normalization and estimated the absorbed and unabsorbed fluxes in $0.3-10\ \rm keV$. We also repeated this analysis while adopting the temperature evolution of SN 2023ixf, $T\propto t^{-0.5}$ (A. J. Nayana et al., 2025), which yielded consistent results to those found with a self-similar solution temperature decline rate. Based on the resulting best‐fit parameters, the X-ray spectra of SN 2024ggi reveal SN ejecta interaction with dense CSM: (i) the emission was highly absorbed at early times, (ii) the absorption decreased with time, and (iii) of thermal nature, consistent with previous studies (R. Margutti et al., 2017; B. P. Thomas et al., 2022). The best-fit parameters are reported in Table 1 and contour plots of the first four epochs data from Swift-XRT and CXO are shown in 4. In addition, the fitted spectrum for CXO and XMM observations are shown in Figure 1.
We classified an observation as a non-detection when the source, SN 2024ggi, has a signal-to-noise ratio below three $(\rm SNR<3)$. For those observations, we used the FakeIt command from PyXspec to simulate fake spectra with the same absorbed thermal Bremsstrahlung model (tbabs*ztbabs*bremss). In the model, we set the intrinsic $\rm N_{H}$ for non-detections to be $\rm N_{H,\ int}=0$ and fixed the temperature to follow $T=34\ \rm keV\ (t/13d)^{-0.25}$. We then used the fake spectrum flux and the count rate ratio to estimate the flux upper-limit in $0.3-10\ \rm keV$ band. However, we did not include parameters derived from non-detections for fitting the luminosity decay and the CSM density profile analysis.
@@
We converted all unabsorbed fluxes to luminosities and fit a power-law with emcee Markov chain Monte Carlo (MCMC). We found that the intrinsic $0.3-10\ \rm keV$ luminosity has a decay of $L_{X}\propto t^{-0.99}$ at $\delta t\gtrsim 8.36$ days. Additionally, in Figure 2, we compare the X-ray light curve to a sample of SNe II compiled by I. Irani et al. (2024); A. J. Nayana et al. (2025).
IV.2 CSM Density Profile Analysis